OpenStudy (anonymous):

Two trains 420 miles apart are headed toward each other. Train A is traveling at 40 mi/hr and Train B at 30 mi/hr. How long will it take before they reach each other? How long will it take before they reach each other? Options are: A-10.5 hours B-14 hours C-4.2 hours D-6 hours

OpenStudy (anonymous):

4.2 Hours

OpenStudy (anonymous):

Thank You :)

OpenStudy (cj49):

Don't give direct answers. >.>

OpenStudy (shinalcantara):

let 't' be the time for both a and b.. velocity A=(distance covered)/t; velocity B=(distance covered)/t; distance of A + distance B =420 equate the two. 40t+30t=420 t=6 hours sweetheart